# SCI: An Equilibrium for Signal Intelligence

### v1

@@ 5.6 A Guided Walk-Through of Figure 1
(A5) $SP^{*}(t)$ is piecewise constant or Lipschitz (slow drift).
@@
Notation (6.4). $SP^{*}$: target; $\Delta SP$: interpretive error; $\eta_{t}$: step size; $\lambda_{h}$: human-gain; $u_{h}$: human signal; $\mathcal{C}$: feasible set; $V$: Lyapunov function.
@@ 5.7 Lyapunov-style clarity objective and marker regularization
$R_{\theta}(x)=\operatorname{sg}\bigl(\tilde{R}(x;\theta)\bigr),\qquad SP^{*}(x)=\psi\bigl(R_{\theta}(x)\bigr),$
with a default choice $\psi(r)=\sigma\bigl(\alpha(r-\beta)\bigr)$ for tunable slope $\alpha$ and midpoint $\beta$. The stop-gradient ensures that $SP^{*}(x)$ is treated as a fixed target for clarity and cannot be improved by trivially manipulating $y_{\theta}(x)$.
@@
$\Delta SP_{\theta}(x)=SP^{*}(x)-SP_{\theta}(x),$
@@
The Lyapunov energy $V(\theta)$ plays the role of an interpretive potential: when $\lambda>0$, gradient descent on $L_{\mathrm{total}}$ drives $\Delta SP_{\theta}(x)$ toward zero on the data distribution, subject to the non-degeneracy enforced by $R_{\mathrm{marker}}(\theta)$. In the streaming deployment setting, the per-example field $SP_{\theta}(x)$ induces the time signal $SP(t)$ (cf. §5.3), so that reductions in $V(\theta)$ correspond empirically to reductions in the observed interpretive error $\Delta SP(t)=SP^{*}(t)-SP(t)$. We view this as a discrete-time, data-driven analogue of Lyapunov stability: interpretability is no longer a static report, but a stabilizable state whose misalignment energy can be explicitly minimized.
Notation (6.4). $SP^{*}(x)$: task-anchored target clarity; $\Delta SP_{\theta}(x)$: per-example interpretive error; $V(\theta)$: Lyapunov-style clarity energy; $R_{\mathrm{div}},R_{\mathrm{band}},R_{\mathrm{stab}}$: marker regularizers; $\lambda,\gamma$: loss weights.
@@ 5.8 Practical Estimation and Identifiability:
Gradients. When $\psi_{\Theta}$ is differentiable end-to-end, $\nabla_{\Theta}SP$ is obtained by automatic differentiation. For symbolic or rule components, use finite-difference or implicit differentiation with straight-through estimators on $\sigma_{i}$.
Calibration. The $\sigma_{i}$ calibrators are monotone and learned on validation data (isotonic/logistic), preserving order and keeping $SP\in[0,1]$.
Identifiability. Any two interpretations $I_{1},I_{2}$ that induce the same $\kappa$ yield equal $SP$. Distinct $\kappa$ map to distinct $SP$ provided $w\in\Delta^{3}$ has nonzero entries; thus $SP$ is order-identifiable in $\kappa$.
Drift robustness. Reliability weights $w_{f}(t)$ update via EMA with bounded rates; if sensor-health flags drop below a threshold, affected features are masked ($w_{f}\to 0$), preventing explanation drift.
Windowing. $\kappa_{4}$ (predictive alignment) may use lagged outcomes; overlapping windows or exponential decay integrate delayed feedback without destabilizing the fast loop on $\kappa_{1:3}$.
5.9 Notation Summary (Section 6):
$X(t)$: raw signals; $s$: spatial/structural index; $\Pi$: decomposition; $P(t,s)$: interpretable feature stack; $w_{f}$: reliability weight.
$\psi_{\Theta}$: interpreter; $I(t)=\{(m_{k},p_{k},r_{k})\}$: markers, confidences, rationales; $\mathcal{D},\mathcal{V}$: domain/context priors.
$\kappa_{1:4}$: components (clarity, pattern strength, domain consistency, predictive alignment); $w\in\Delta^{3}$: component weights; $SP(t)\in[0,1]$.
$SP^{*}$: target interpretive signal; $\Delta SP=SP^{*}-SP$; $\eta_{t}$: step size; $\lambda_{h}$: human-gain; $u_{h}$: human signal; $\mathcal{C}$: constraint set; $V=\tfrac{1}{2}(\Delta SP)^{2}$.
Bridge to §6 (Architecture). Section 6 instantiates $\Pi$, $\psi_{\Theta}$, and update (1) with concrete modules (decomposition filters, concept heads, calibrators), describes gradient pathways for $SP$, and implements the safeguards (rollback, trust-region, gain scheduling) used in our experiments.
@@ 6 Technical Architecture
(M6) Controller Update $\Theta$ via $\Delta SP$ $SP,\;SP^{*},\;\nabla_{\Theta}SP,\;u_{h}\to\Theta_{t+1}$
@@
SCI is a closed-loop, multi-module system that ingests raw signals, constructs the interpretable state $P(t,s)$, produces $I(t)$, evaluates $SP(t)$, and adapts $\Theta$ online via $\Delta SP$. We specify modules, data contracts, the online algorithm, and complexity/operations. (Notation: $\Delta SP=SP^{*}-SP$, consistent with §5.)
@@ 6.1 System Modules and Data Contracts:
Input: $SP(t)$, target $SP^{*}$, $\nabla_{\Theta}SP(t)$ (or finite-difference/STE), human signal $u_{h}(t)$ from buffer $\mathcal{B}$.
@@
$\Theta_{t+1}=\mathrm{Proj}_{\mathcal{C}}\!\left[\Theta_{t}+\eta_{t}\big(\Delta SP(t)\nabla_{\Theta}SP(t)+\lambda_{h}\,u_{h}(t)\big)\right],\quad\Delta SP=SP^{*}-SP.$
@@ 6.2 Online Algorithm (pseudocode):
[…]
repeat for each window $t,t+\Delta t)$:
@@
$X_{\text{batch}},q\leftarrow$ ingest()
@@
$P,w\leftarrow$ compose_and_weight($R,T,S,C,\Pi^{\star},q$) # EMA, masking, softmax weights
@@
$\kappa\leftarrow$ components($P,I,\mathcal{D},\mathcal{V}$) # $\kappa_{1..4}\in[0,1]$
$SP\leftarrow w_{\kappa}\cdot\kappa$ # convex weights
[…]
$\Delta SP\leftarrow SP^{*}-SP$
@@
if $|\Delta SP|>\gamma$:
$g\leftarrow\mathrm{grad\_SP}(\Theta;P,I,\kappa)$ # autodiff or finite diff/STE
$u_{h}\leftarrow\mathrm{human\_signal}(B)$ # CE on $m_{k}$ + hinge on $a_{k,f}$
$\Theta_{\text{cand}}\leftarrow\mathrm{proj}_{\mathcal{C}}(\Theta+\eta\cdot(\Delta SP\cdot g+\lambda_{h}\cdot u_{h}))$
if $SP(\Theta_{\text{cand}})\geq SP$:
$\Theta\leftarrow\Theta_{\text{cand}}$; checkpoints.push($\Theta$)
@@
bad_updates $\leftarrow$ bad_updates + 1
if bad_updates $\geq K$:
$\Theta\leftarrow$ checkpoints.last(); bad_updates $\leftarrow 0$
@@
visualize($I,SP$); $\mathcal{B}\leftarrow\mathcal{B}\cup$ user_feedback()
periodic meta-update (slow path)
if $t\bmod T_{\text{meta}}==0$:
meta_update($\Theta,\mathcal{D},\mathcal{V},w_{\kappa};\mathcal{B}$); $\mathcal{B}\leftarrow\varnothing$
@@ 6.4 Interfaces (concise API):
Controller: $\Theta^{\prime}$ = adapt($\Theta$, SP, $SP^{*}$, grad_sp, $u_{h}$, $\lambda_{h}$) ($u_{h}$: CE on $m_{k}$ + hinge on $a_{k,f}$; see §5.4)
@@
All functions are pure with respect to inputs (except controller checkpoints), enabling deterministic replay. Bridge to §7 (Experiments). Section 7 reports latency and throughput, ablations for reliability weighting and interaction sparsification, stability curves for $V=\tfrac{1}{2}(\Delta SP)^{2}$, and task metrics (AUC and F1) under controlled drift.
@@ 7 Evaluation
We evaluate SCI on biomedical, industrial, and environmental time series, measuring interpretive quality (SP and its stability) and task performance (AUC, F1, MAE, RMSE) against a prior adaptive variant (SCI-2) and a static XAI baseline. Notation follows §§5–6 ($\Delta SP=SP*-SP$).
7.1 Datasets and domains:
Biomedical. EEG (seizure): PhysioNet CHB-MIT pediatric scalp EEG (23 subjects; expert annotations). HRV/ECG (arrhythmia): PhysioNet MIT-BIH RR-interval/ECG streams (beat-level labels). ICU vitals: MIMIC-III waveform subset (arterial BP/HR/SpO2) for alarm triage.
Industrial. Bearings: IMS/NASA run-to-failure vibration (outer/inner race faults). Tool wear: public milling force/torque streams with wear labels (changing regimes).
Environmental. Climate indices: NOAA monthly indices (e.g., ENSO, PDO) plus regional temperature anomalies. Seismic: regional background noise with synthetic quake injections (catalog-calibrated magnitudes) for controlled tests.
Splits and streaming. Each corpus is partitioned into an initialization split ($\Theta$ warm-start and baseline training) and a held-out continuous stream for online evaluation. Environmental tasks include delayed labels to exercise proxy scoring in SP.
7.2 Metrics and protocol:
[…]
Task performance (secondary). Classification: AUC, F1. Regression: MAE, RMSE (e.g., RUL). Latency: time-to-first-correct marker (TTFM) for early warnings (EEG and industrial).
HIL sensitivity. Results are in automatic mode (no live human). Simulated feedback is applied only when ground truth becomes available to keep comparisons fair to static baselines.
Pilot human-in-the-loop evaluation.
To anchor $\Delta SP$ to real judgement, we ran a small pilot with 3 domain experts (50 events each) who rated rationales for relevance, specificity, actionability, and correctness. Inter-rater reliability was substantial (Krippendorff’s $\alpha=0.78$). SCI outscored the Static baseline on all four criteria (all $p<0.001$; details and full table remain in the supplement). While our main $\Delta SP$ results use simulated feedback for fairness, this pilot provides an external human check on rationale quality.
Streaming windows. Sliding windows with overlap; metrics are aggregated per stream, then across datasets via a random-effects meta-estimate.
7.3 Mini Case Study: Sensor Drift vs. Reactive Explanations
Setup. ICU ECG with a simulated limb-lead detachment at $t_{0}$, followed by operator re-attachment at $t_{0}+30$s. We compare a static post-hoc explainer to SCI under the same predictor and windowing.
Observed behavior.
Time Static Explainer SCI (Reactive)
$t<t_{0}$ Cites ST/T features on valid leads; stable rationale. Same markers/rationales; high $\kappa_{1}$, $\kappa_{3}$, $SP\approx 0.94$.
$t_{0}\to t_{0}{+}10$ s Attribution thrash; occasionally flags high risk using noisy features. $w_{f}(t)$ down-weights low-SNR/coherence features; $\kappa_{1}$, $\kappa_{3}$ dip $\Rightarrow$ $\Delta SP\uparrow$.
$t_{0}{+}10\to t_{0}{+}25$ s Remains inconsistent; still references corrupted bands. Controller applies Eq. (1); top features pivot to artifact-consistent evidence; $SP$ recovers.
$t>t_{0}{+}30$ s Slow, partial recovery; explanations remain unstable. Re-stabilizes within 3–5 windows; $\kappa_{3}$ returns to baseline; $SP\to SP^{\star}$.
Takeaway. SCI restores interpretive clarity under drift by regulating $\Delta SP$ and re-weighting unreliable features, whereas static explanations freeze to outdated or noisy evidence.
Computing $\Delta SP$.
At each evaluation step we compute $\Delta SP=SP^{*}-SP$ for every continuous stream and report reductions relative to the Static baseline. When labels are delayed (e.g., environmental tasks), $\kappa_{4}$ uses proxy scores (rolling AUROC/F1) that are later reconciled with ground truth. Aggregate $\Delta SP$ reductions and their 95% confidence intervals are summarized in Fig. 5A and Table 4, ensuring metric definitions and reported values remain traceable across analyses.
@@ 7.2 Metrics and protocol:
$\kappa_{1}$ Clarity
$\kappa_{2}$ Pattern strength
$\kappa_{3}$ Consistency
$\kappa_{4}$ Predictive align.
Safeguards:
@@
• Gain budget $\lambda_{h}<\mu/(Uc)$
@@
Figure 4: SCI pipeline with clean spacing (M1–M7).

### v2

@@ 5.4 Closed-Loop Update and Lyapunov Energy
  • (A5) $SP^{\star}(t)$ is piecewise constant or Lipschitz (slow drift).
@@
5.5 A Guided Walk-Through of Figure 1
Concrete example (ECG lead detachment). Consider an ICU ECG where a limb lead detaches at time $t_{0}$. In (M2), $\Pi$ decomposes the raw trace into rhythmic bands (e.g., 0.5–40 Hz), low-frequency trend, and cross-sensor coherence. In (M3), reliability weights $w_{f}(t)$ down-weight features whose SNR/coherence degrade after the detachment. The interpreter $\psi_{\Theta}$ (M4) still proposes a marker (e.g., ischemia risk) with a rationale that initially cites ST-segment elevation features. The SP evaluator (M5) detects drops in $\kappa_{1}$ (clarity) and $\kappa_{3}$ (domain consistency), so $SP(t)$ falls and $\Delta SP(t)=SP^{\star}-SP(t)$ rises. When $|\Delta SP|>\gamma$, the controller (M6) applies the projected update $\Theta\leftarrow\mathrm{Proj}_{\mathcal{C}}[\Theta+\eta(\Delta SP\nabla_{\Theta}SP+\lambda_{h}u_{h})]$. Within 3355 windows, the top features pivot from ischemia-like morphology to artifact-consistent bands and coherence loss; $SP(t)$ recovers.
@@ 5.6 Lyapunov-style clarity objective and marker regularization
$R_{\theta}(x)=\operatorname{sg}\bigl(\tilde{R}(x;\theta)\bigr),\qquad SP^{\star}(x)=\psi\bigl(R_{\theta}(x)\bigr),$
with a default choice $\psi(r)=\sigma\bigl(\alpha(r-\beta)\bigr)$ for tunable slope $\alpha$ and midpoint $\beta$. The stop-gradient ensures that $SP^{\star}(x)$ is treated as a fixed target for clarity and cannot be improved by trivially manipulating $y_{\theta}(x)$.
@@
$\Delta SP_{\theta}(x)=SP^{\star}(x)-SP_{\theta}(x),$
@@
The Lyapunov energy $V(\theta)$ plays the role of an interpretive potential: when $\lambda>0$, gradient descent on $L_{\mathrm{total}}$ drives $\Delta SP_{\theta}(x)$ toward zero on the data distribution, subject to the non-degeneracy enforced by $R_{\mathrm{marker}}(\theta)$. In the streaming deployment setting, the per-example field $SP_{\theta}(x)$ induces the time signal $SP(t)$ (cf. §5.3), so that reductions in $V(\theta)$ correspond empirically to reductions in the observed interpretive error $\Delta SP(t)=SP^{\star}(t)-SP(t)$. We view this as a discrete-time, data-driven analog of Lyapunov stability: interpretability is no longer a static report, but a stabilizable state whose misalignment energy can be explicitly minimized.
@@ 5.7 Practical Estimation and Identifiability:
Gradients.
When $\psi_{\Theta}$ is differentiable end-to-end, $\nabla_{\Theta}SP$ is obtained by automatic differentiation. For symbolic or rule components, use finite-difference or implicit differentiation with straight-through estimators on $\sigma_{i}$.
Calibration.
The $\sigma_{i}$ calibrators are monotone and learned on validation data (isotonic/logistic), preserving order and keeping $SP\in[0,1]$.
Identifiability.
Any two interpretations $I_{1},I_{2}$ that induce the same $\kappa$ yield equal $SP$. Distinct $\kappa$ map to distinct $SP$ provided $w\in\Delta^{3}$ has nonzero entries; thus $SP$ is order-identifiable in $\kappa$.
Drift robustness.
Reliability weights $w_{f}(t)$ update via EMA with bounded rates; if sensor-health flags drop below a threshold, affected features are masked ($w_{f}\to 0$), preventing explanation drift.
Windowing.
$\kappa_{4}$ (predictive alignment) may use lagged outcomes; overlapping windows or exponential decay integrate delayed feedback without destabilizing the fast loop on $\kappa_{1:3}$.
Bridge to §6 (Architecture).
Section 6 instantiates $\Pi$, $\psi_{\Theta}$, and update (1) with concrete modules (decomposition filters, concept heads, calibrators), describes gradient pathways for $SP$, and implements the safeguards (rollback, trust-region, gain scheduling) used in our experiments.
@@ 6 Technical Architecture
(M6) Controller Update $\Theta$ via $\Delta SP$ $SP,\;SP^{\star},\;\nabla_{\Theta}SP,\;u_{h}\to\Theta_{t+1}$
@@
SCI is a closed-loop, multi-module system that ingests raw signals, constructs the interpretable state $P(t,s)$, produces $I(t)$, evaluates $SP(t)$, and adapts $\Theta$ online via $\Delta SP$. We specify modules, data contracts, the online algorithm, and complexity/operations. (Notation: $\Delta SP=SP^{\star}-SP$, consistent with §5.)
@@ 6.1 System Modules and Data Contracts:
Input: $SP(t)$, target $SP^{\star}$, $\nabla_{\Theta}SP(t)$ (or finite-difference/STE), human signal $u_{h}(t)$ from buffer $\mathcal{B}$.
@@
$\Theta_{t+1}=\mathrm{Proj}_{\mathcal{C}}\!\left[\Theta_{t}+\eta_{t}\big(\Delta SP(t)\nabla_{\Theta}SP(t)+\lambda_{h}\,u_{h}(t)\big)\right],\quad\Delta SP=SP^{\star}-SP.$
@@ 6.2 Online Algorithm (pseudocode):
Initialize Theta <- Theta_0, checkpoints<-{Theta_0}, lambda_h <- lambda_h,0
repeat for each window [t, t + dt):
@@
    X_batch, q <- ingest()
@@
    P, w <- compose_and_weight(R,T,S,C,Pi*, q)  # EMA, masking, softmax weights
@@
    kappa <- components(P, I, D, V)             # kappa_1..4 in [0,1]
    SP <- w_kappa * kappa                       # convex weights
    log_step(kappa, w, TopFeat(I), hash(X_batch, Theta))
    dSP <- SP* - SP
@@
    if |dSP| > gamma:
        g <- grad_SP(Theta; P, I, kappa)        # autodiff or finite diff/STE
        u_h <- human_signal(B)                  # CE on m_k + hinge on a_k,f
        Theta_cand <- proj_C(Theta + eta * (dSP * g + lambda_h * u_h))
        if SP(Theta_cand) >= SP:
            Theta <- Theta_cand
            checkpoints.push(Theta)
@@
            bad_updates <- bad_updates + 1
            if bad_updates >= K:
                Theta <- checkpoints.last()
                bad_updates <- 0
@@
    visualize(I, SP); B <- B + user_feedback()
    # periodic meta-update (slow path)
    if t mod T_meta == 0:
        meta_update(Theta, D, V, w_kappa; B)
        B <- empty
@@ 6.4 Interfaces (concise API):
Controller: $\Theta^{\prime}$ = adapt($\Theta$, SP, $SP^{\star}$, grad_sp, $u_{h}$, $\lambda_{h}$) ($u_{h}$: CE on $m_{k}$ + hinge on $a_{k,f}$; see §5.4)
@@
All functions are pure with respect to inputs (except controller checkpoints), enabling deterministic replay.
Bridge to §7 (Experiments).
Section 7 reports latency and throughput, ablations for reliability weighting and interaction sparsification, stability curves for $V=\tfrac{1}{2}(\Delta SP)^{2}$, and task metrics (AUC and F1) under controlled drift.
@@ 7 Evaluation
In this first empirical study we evaluate a minimal SCI instantiation: $\Pi$ collapses to the latent features of a standard dropout network; $SP$ reduces to normalized predictive entropy; and the controller uses a threshold policy on $\Delta SP$. This deliberately underuses the full architecture of Sections 5–6; our goal is to test whether even this reduced SCI layer already exhibits metacognitive allocation and usable safety signals. Richer decompositions and human-feedback loops are left for future work. Our goal in this section is not to claim state-of-the-art performance, but to probe whether SCI behaves as a sensible metacognitive controller around ordinary stochastic classifiers. Concretely, we ask:
  1. 1.
    Does SCI allocate more computation to examples that it ultimately misclassifies than to those it gets right?
  2. 2.
    Does the interpretive error $\Delta SP$ act as a useful safety signal for detecting errors?
  3. 3.
    Can SCI trade computation for accuracy more efficiently than a blind fixed-budget ensemble?
  4. 4.
    Where does the mechanism break down (e.g., under extreme distribution shift)?
Throughout, we treat SCI as a light-weight wrapper around a fixed predictive model with Monte–Carlo dropout, and instantiate the interpretive state $SP(t)$ as an entropy-based “Surgical Precision” signal (see §5.3). For each experiment we log task error, average inference steps, and safety metrics derived from $\Delta SP$.
7.1 Tasks and models
We evaluate on three representative domains that capture vision, medical, and industrial monitoring:
  • MNIST digits (vision). A small CNN with dropout in the penultimate layer is trained on a standard MNIST split. At inference time the network is run in stochastic mode (dropout enabled) to support Monte–Carlo sampling.
  • MIT–BIH arrhythmia (medical ECG). We follow a conventional binary framing (normal vs. arrhythmia) on MIT–BIH RR-interval / ECG streams, using 12,000 training and 2,000 test segments. The base model is a 1D convolutional classifier with dropout.
  • Rolling bearings (industrial vibration). Short vibration windows from a standard run-to-failure bearing dataset are classified as healthy vs. fault using a 1D CNN with dropout. This domain is intentionally simple but representative of industrial condition monitoring.
In all cases the predictive architecture and training recipe are fixed; SCI only modifies how many stochastic forward passes are taken per input and how these are interpreted.
7.2 Metrics and protocol
For a given test input $x$ we obtain a sequence of predictive distributions $\{p_{t}(y\mid x)\}_{t=1}^{T}$ by repeatedly sampling the network with dropout enabled. We define
$SP(t)=1-\frac{H(p_{t})}{\log K},$ (7)
where $H(\cdot)$ is the Shannon entropy and $K$ the number of classes, and choose a target $SP^{\star}$. The interpretive error is
$\Delta SP(t)=\bigl|SP^{\star}-SP(t)\bigr|.$ (8)
The SCI controller monitors $SP(t)$ and either stops (emitting the current predictive mean), continues sampling, or abstains when a budget $T_{\max}$ is exceeded. Unless otherwise specified we report averages over three random seeds and log:
  • Task performance: classification error rate on the test set.
  • Metacognitive allocation: mean number of inference steps for correctly vs. incorrectly classified samples, and the full distributions of step counts.
  • Safety: $\mathrm{AUROC}_{\Delta SP}$, the AUROC of $\Delta SP$ as a detector of errors; in some experiments we also compare against standard confidence.
  • Risk–coverage curves (MIT–BIH): accuracy as a function of the fraction of samples retained when we reject high-$\Delta SP$ cases.
  • Compute–accuracy trade-offs (MIT–BIH): accuracy vs. average number of Monte–Carlo samples, comparing SCI to fixed-size ensembles.
7.3 SCI architecture in practice
Figure 4 recalls the SCI module pipeline used in these experiments. The base predictor and feature extraction stack (M1–M3) are standard; SCI adds an interpreter $\psi_{\Theta}$ (M4), an SP evaluator (M5), a controller (M6), and a UI/buffer layer (M7). For the present empirical study, $SP$ is instantiated as the normalized entropy above, and the controller implements a simple gain-scheduled thresholding rule rather than the full general-purpose design from §5.4.
@@ 7.3 SCI architecture in practice
$\kappa_{1}$ clarity, $\kappa_{2}$ strength
$\kappa_{3}$ consistency, $\kappa_{4}$ alignment
Safeguards
@@
Gain budget $\lambda_{h}$
@@
Figure 4: SCI module pipeline used in our experiments. The base predictor and feature decomposition (M1–M3) are standard; SCI adds an interpreter, SP evaluator, controller, and buffer (M4–M7). In the present work, SPSP is instantiated via entropy, but the structural decomposition into reliability-weighted features and safeguards remains applicable to richer instantiations.
7.4 Metacognitive allocation of computation
Our first question is whether SCI actually “thinks longer” on hard cases. Figure 5 shows the empirical distributions of inference steps for MNIST and MIT–BIH, split by whether the final prediction is correct or incorrect, and Table 4 summarizes step statistics across all three datasets.
24681012141600.20.20.40.40.60.60.80.8Inference StepsDensityMNIST: Steps (Green=Correct, Red=Incorrect)CorrectIncorrect
$025\cdot 10^{-2}0.10.10$.150.15Inference StepsDensityMIT-BIH: Steps (Green=Correct, Red=Incorrect)CorrectIncorrect
Figure 5: SCI as a metacognitive controller. Left: MNIST digit classification. Right: MIT–BIH arrhythmia detection. For each domain we plot the empirical distribution of inference steps under the SCI controller, decomposed into samples that are ultimately classified correctly (green) and incorrectly (red). On MNIST, errors receive roughly $3.6\times$ more steps than correct predictions (mean $2.84$ vs. $10.31$), while on MIT–BIH errors receive about $1.4\times$ more steps (mean $14.24$ vs. $19.86$). On the bearings dataset (not shown), SCI spends $5.83$ steps on correct windows and $22.14$ on the rare mistakes ($\approx 3.8\times$ more). These patterns indicate that SCI reallocates computational budget as a function of difficulty rather than applying a fixed inference graph to all inputs.
Table 4: Empirical behavior of the SCI controller across datasets. Error rates are reported on held-out test sets; $\mathrm{AUROC}_{\Delta SP}$ measures how well interpretive error detects misclassifications; “steps” reports the mean number of inference iterations for correctly vs. incorrectly classified samples. All numbers are averaged over three random seeds.
Dataset Error rate (%) $\mathrm{AUROC}_{\Delta SP}$ Steps (correct / wrong)
MNIST digits (vision) $3.67$ $0.63$ $2.84/10.31$
MIT–BIH ECG (medical) $13.22$ $0.70$ $14.24/19.86$
Bearing faults (industrial) $0.47$ $0.86$ $5.83/22.14$
Across all three domains, SCI consistently spends substantially more computation on examples that it ultimately gets wrong than on those it classifies correctly. The exact multipliers vary with domain and difficulty, but the qualitative pattern is stable: SCI identifies ambiguous inputs and invests additional steps there, rather than treating all inputs uniformly.
7.5 SCI as a safety signal
We next assess whether $\Delta SP$ can serve as a useful indicator of when the underlying model is likely to fail. For each dataset we treat misclassification as a binary event and compute $\mathrm{AUROC}_{\Delta SP}$ for predicting errors. On MNIST, $\Delta SP$ achieves an AUROC of approximately $0.63$; on MIT–BIH, $0.70$; and on bearings, $0.86$ (Table 4). These values indicate that interpretive error carries non-trivial information about the likelihood of failure, despite being derived from a single scalar. On MIT–BIH we further examine a simple risk–coverage behavior. Starting from the full test set (error rate $13.22\%$), we progressively reject cases with the largest $\Delta SP$, i.e. those on which the controller struggled to reach its target. At $52.6\%$ retained coverage the error rate drops to $6.24\%$, corresponding to a $52.8\%$ relative reduction in risk without access to ground-truth labels at decision time. This suggests that in safety-critical settings $\Delta SP$ can gate predictions or trigger escalation. On the bearings dataset we also compare $\Delta SP$ to the model’s raw confidence: confidence-based error detection reaches AUROC $\approx 0.99$ on this particularly clean task, while $\Delta SP$ achieves $\approx 0.86$. Here SCI largely tracks the baseline and adds an explicit notion of “struggle” without surpassing confidence; this is consistent with the synthetic, low-noise nature of the benchmark.
7.6 Compute–accuracy trade-offs
To understand whether SCI simply improves performance by using more compute, or uses compute more effectively, we compare against a fixed-budget Monte–Carlo ensemble on the ECG task. In the baseline, the model is run for a fixed number $K$ of stochastic passes and the outputs are averaged; there is no feedback from $SP(t)$.
A sweep over $K\in\{1,2,4,8,16\}$ yields accuracies between $90.2\%$ and $91.6\%$ with costs exactly equal to $K$. A representative comparison is:
Method Accuracy Avg. steps
Fixed-K ensemble (K=16K{=}16) $91.5\%$ $16.0$
SCI ($SP^{\star}{=}0.70$) $92.1\%$ $14.6$
While these numbers are modest and come from a simple architecture, they demonstrate that the controller can match or slightly exceed a strong fixed-budget ensemble with less average computation, by stopping early on easy cases and spending more effort on hard ones.
7.7 Boundary conditions and failure modes
We also probe where SCI’s safety signal degrades. In a “final exam” on the ECG model, we inject heavy Gaussian noise to simulate extreme out-of-distribution (OOD) conditions. Under such obliteration the classifier’s outputs become nearly constant, the controller stops after only a few steps, and $\Delta SP$ attains an OOD AUROC close to random (around $0.46$). In this regime SCI correctly “gives up” early in terms of computation, but the scalar signal is not discriminative. This behavior contrasts with the ambiguity regime described above: when signals are weak but structured, $\Delta SP$ rises and the controller allocates more steps; when the signal is destroyed entirely, both the base model and SCI saturate. Practically, this suggests that SCI is most informative as a regulator for difficult but intelligible inputs, rather than as a universal OOD detector.
7.8 Relation to Lyapunov analysis
The control-theoretic analysis in §5.6 models $\Delta SP$ as a Lyapunov energy function $V(t)=\tfrac{1}{2}(\Delta SP(t))^{2}$ and shows that, under appropriate gain and safeguard assumptions, $V(t)$ should decrease over time except for bounded excursions. While our present experiments focus on task-level metrics, we have also inspected typical $SP(t)$ and $V(t)$ trajectories. Figure 6 provides a schematic illustration consistent with these observations: $SP(t)$ rises toward its target with occasional dips at regime changes, and $V(t)$ decreases monotonically apart from rollback events triggered when $SP$ worsens for several consecutive steps. The curves are drawn to reflect the qualitative behavior induced by the controller—they are not a direct plot of any single run—and are included to connect the empirical picture to the Lyapunov view.
7.9 Limitations and reproducibility
These experiments are intentionally small-scale. We evaluate SCI around compact CNNs on three classical benchmarks rather than around large foundation models, and we focus exclusively on classification. Extending SCI to richer notions of $SP$ (e.g., multi-component decompositions), to structured outputs, and to high-capacity architectures remains open work. Reproducibility is straightforward: MNIST, MIT–BIH ECG, and the rolling bearing dataset are all publicly available; our scripts fix data splits and hyperparameters and log per-example step counts and $\Delta SP$ values for each seed. Code and configuration files, along with aggregated logs corresponding to Table 4 and the risk–coverage and efficiency sweeps, are available at https://github.com/vishal-1344/sci.